6.17 Suppose that amplifiers have a constant failure rate of λ=0.8/month. Suppose that four such amplifiers are tested for 6 months. What is the probability that more than one of them will fail? Assume that when they fail, they are not replaced.
Answer #1

Here, λ = 4.8 and x = 1
As per Poisson's distribution formula P(X = x) = λ^x * e^(-λ)/x!

We need to calculate P(X > 1) = 1 - P(X <= 1).
P(X > 1) = 1 - (4.8^0 * e^-4.8/0!) + (4.8^1 * e^-4.8/1!)
P(X > 1) = 1 - (0.0082 + 0.0395)
P(X > 1) = 1 - 0.0477
= 0.9523

Ans: 0.9523